9 defective electric toothbrushes were accidentally shipped to a drugstore by Cleanbrush Products along with 29 nondefective ones.

What is the probability the first two electric toothbrushes sold will not be defective?


1
Expert's answer
2022-05-31T11:17:07-0400

Let "D" denote the event "electric toothbrush is defective".


"P(D^CD^C)=\dfrac{29}{9+29}(\dfrac{28}{9+28})=\dfrac{406}{703}\approx0.5775"